a. When a solid dissolves in water, heat may be evolved or absorbed. The heat of dissolution (dissolving) can be determined using a coffee cup calorimeter.

In the laboratory a general chemistry student finds that when 16.11 g of BaCl2(s) are dissolved in 118.70 g of water, the temperature of the solution increases from 22.98 to 25.41 °C.

The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.71 J/°C.

Based on the student's observation, calculate the enthalpy of dissolution of BaCl2(s) in kJ/mol.

Assume the specific heat of the solution is equal to the specific heat of water.

ΔHdissolution =  kJ/mol

b.When a solid dissolves in water, heat may be evolved or absorbed. The heat of dissolution (dissolving) can be determined using a coffee cup calorimeter.

In the laboratory a general chemistry student finds that when 3.22 g of CuSO4(s) are dissolved in 110.60 g of water, the temperature of the solution increases from 24.18 to 27.36 °C.

The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.64 J/°C.

Based on the student's observation, calculate the enthalpy of dissolution of CuSO4(s) in kJ/mol.

Assume the specific heat of the solution is equal to the specific heat of water.

ΔHdissolution =  kJ/mol
Answer #1

Mass of BaCl2 = 16.77 g Number of moles of BaCl, = mass/mol.wt. = 16.77 g/(208.23 g/mol) = 0.080536 mol Mass of water = 118.7

==========================================================================================================

Mass of CuSO4 = 3.22 g Number of moles of CuSO4 = mass/mol.wt. = 3.22 g/(159.61 g/mol) = 0.020174 mol Mass of water = 110.60